OpenStudy (anonymous):

Daniel set a goal for the number of servings of fruits and vegetables he wants to eat each day. He has already eaten four and one-fifth servings today, and he still needs to eat three and three-fifths servings to meet his goal. What is his goal for daily servings of fruits and vegetables? one and two-fifths six and three-fifths seven and two-fifths seven and four-fifths

OpenStudy (anonymous):

help pleas

OpenStudy (anonymous):

can you help pleas

OpenStudy (anonymous):

\[4\frac{ 1 }{ 5 }+3\frac{ 3 }{ 5 }\]becomes\[\frac{ 21 }{ 5 }+\frac{ 18 }{ 5 }\]Do you know how to add fractions?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok, so you need a common denominator, which you already do since they are both 5. So you can add 21+18. What's 21+18?

OpenStudy (anonymous):

Or here. This will be easier.\[4\frac{ 1 }{ 5 }+3\frac{ 3 }{ 5 }\]Since you have the same denominator you can ad 4 and 3 to give you 7. then add 1 and 3. That will give you 4. Do you know what the answer is?